Find the height of a trapezoid with area of 99m2

 and the bottom base is twice the top base, 

which measures 6m.

1
Expert's answer
2021-12-16T09:58:07-0500

Calculation


  • You can use the equation for the are of a trapezoid which is "\small A = \frac{1}{2}(l_1+l_2)h" . Here "\small l" represents the length of any base.
  • If the bottom base is "\small 6\,m" then the top base is "\small 3\,m" .
  • Upon substitution the values,

"\qquad\qquad
\begin{aligned}
\small 99\,m^2&=\small \frac{1}{2}(6\,m+3\,m)h\\
\small 99\times2&=\small \cancel{2}\times\frac{1}{\cancel{2}}\times(9)\times h\\
\small 99\times2\times\frac{1}{9}&=\small\frac{1}{\cancel{9}}\times \cancel{9}h\\
\small h &=\small 2
\end{aligned}"